A nutritionist wants to know the population of Kindergarten pupils who eat

vegetables. A survey among 1500 kinder pupils was conducted and revealed

that 485 pupils ate vegetables. Construct the 99% confidence interval to

determine the true proportion of pupils who eat vegetables. Interpret the data.


1
Expert's answer
2021-10-19T12:30:39-0400

The sample proportion is computed as follows, based on the sample size "N = 1500" and the number of favorable cases "X = 485:"


"\hat{p}=\dfrac{485}{1500}=\dfrac{97}{300}\approx0.3233"

The critical value for "\alpha = 0.01" is "z_c = z_{1-\alpha/2} = 2.5758."

The corresponding confidence interval is computed as shown below:


"CI(Proporttion)=(\hat{p}-z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}},"

"\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})"

"=(\dfrac{97}{300}-2.5758\sqrt{\dfrac{\dfrac{97}{300}(1-\dfrac{97}{300})}{1500}},"

"\dfrac{97}{300}+2.5758\sqrt{\dfrac{\dfrac{97}{300}(1-\dfrac{97}{300})}{1500}})"

"=(0.2922,0.3544)"

Therefore, based on the data provided, the 99% confidence interval for the population proportion is "0.2922 < p < 0.3544," which indicates that we are 99% confident that the true population proportion "p"  is contained by the interval "(0.2922, 0.3544)."